Full Abstraction of PCF and Related Languages

A Collection of Literature

December 1, 1995

This document is also available as PostScript, DVI, Text.

References

[BE94] Abstract, PostScript, DVI.

Antonio Bucciardelli and Thomas Ehrhard. Sequentiality in an extensional framework. Information and Computation, 110:265–296, 1994.

[BG92] Abstract, PostScript.

Steven Brookes and Shai Geva. Stable and sequential functions on Scott domains. June 1992. Submitted for publication.

[BG93] Abstract, PostScript.

Steven Brookes and Shai Geva. Sequential functions on indexed domains and full abstraction for a sub-language of PCF. 1993. In Mathematical Foundations of Programming Semantics, Proceedings of the Ninth International Conference, MFPS ’93, New Orleans, April 1993. Lecture Notes in Computer Science, vol 802. Springer-Verlag.

[Ehr93] Abstract, PostScript, DVI.

Thomas Ehrhard. Hypercoherences: a strongly stable model of linear logic. MSCS - Mathematical Structures in Computer Science, 2:365–385, 1993. Slightly changed version.

[Ehr94] Abstract, PostScript, DVI.

Thomas Ehrhard. Projecting sequential algorithms on strongly stable functions. 44 pp. To appear in Annals of Pure and Applied Logic, 1994.

[HO94] Abstract, PostScript.

J. Martin E. Hyland and C.-H. Luke Ong. On full abstraction for PCF: I, II and III (preliminary version). 109 pp, 1994.

[HO95] Abstract, PostScript.

J. Martin E. Hyland and C.-H. Luke Ong. Pi-calculus, dialogue games and PCF. In Proceedings of 7th Annual ACM Conference on Functional Programming Languages and Computer Architecture (La Jolla, California, June 26–28, 1995). ACM - Association for Computing Machinery, 1995.

[Sto88] Abstract, PostScript.

Allen Stoughton. Fully Abstract Models of Programming Languages, chapter Preface, Introduction, Bibliography and Index, pages i–iv, 1–7 and 119–123. Research Notes in Theoretical Computer Science. Pitman, 1988.

[Sto90] Abstract, PostScript. Allen Stoughton. Equationally fully abstract models of PCF. In Michael Main, Austin Melton, Michael Mislove, and David Schmidt, editors, Mathematical Foundations of Programming Semantics: 5th International Conference (Tulane University, New Orleans, Louisiana, 1989, March/April), number 442 in Lecture Notes in Computer Science, pages 271–283, Berlin, 1990. Springer-Verlag.

[Sto91a] Abstract, PostScript. Allen Stoughton. Interdefinability of parallel operations in PCF. Theoretical Computer Science - Journal of EATCS, 79:357–358, 1991.

[Sto91b] Abstract, PostScript. Allen Stoughton. Parallel PCF has a unique extensional model. In Proceedings, Sixth Annual IEEE Symposium on Logic in Computer Science, pages 146–151, Amsterdam, The Netherlands, July 15–18 1991. IEEE Computer Society Press.

[Sto93] Abstract, PostScript. Allen Stoughton. Studying the fully abstract model of PCF within its continuous function model. In Marc Bezem and Jan Friso Groote, editors, Typed Lambda Calculi and Applications: International Conference on Typed Lambda Calculi and Applications, TLCA ’93 (Utrecht, The Netherlands, 1993, March 16-18), number 664 in Lecture Notes in Computer Science, pages 230–244, Berlin, 1993. Springer-Verlag.

[Sto94] Abstract, PostScript. Allen Stoughton. Mechanizing logical relations. In Stephen Brookes et al., editors, Mathematical Foundations of Programming Semantics: 9th International Conference proceedings (New Orleans, LA, USA, 1993, April 7–10), number 802 in Lecture Notes in Computer Science, pages 359–377, Berlin, 1994. Springer-Verlag. Corrected version.

[Sto] HTML.

Allen Stoughton. Lambda.